### v1

# An efficient wavelet-based physics-informed neural networks for singularly perturbed problems

Abstract
Physics-informed neural networks (PINNs) are a class of deep learning models that utilize physics in the form of differential equations to address complex problems, including ones that may involve limited data availability. However, tackling solutions of differential equations with oscillations or singular perturbations and shock-like structures becomes challenging for PINNs. Considering these challenges, we designed an efficient wavelet-based PINNs (W-PINNs) model to solve singularly perturbed differential equations. Here, we represent the solution in wavelet space using a family of smooth-compactly supported wavelets. This framework represents the solution of a differential equation with significantly fewer degrees of freedom while still retaining in capturing, identifying, and analyzing the local structure of complex physical phenomena. The architecture allows the training process to search for a solution within wavelet space, making the process faster and more accurate. Further, the proposed model does not rely on automatic differentiations for derivatives involved in differential equations and does not require any prior information regarding the behavior of the solution, such as the location of abrupt features. Thus, through a strategic fusion of wavelets with PINNs, W-PINNs excel at capturing localized nonlinear information, making them well-suited for problems showing abrupt behavior in certain regions, such as singularly perturbed problems. The efficiency and accuracy of the proposed neural network model are demonstrated in various test problems, i.e., highly singularly perturbed nonlinear differential equations, the FitzHugh-Nagumo (FHN), and Predator-prey interaction models. The proposed design model exhibits impressive comparisons with traditional PINNs and the recently developed wavelet-based PINNs, which use wavelets as an activation function for solving nonlinear differential equations.
keywords:
Physics-informed neural networks, Wavelets, Singularly perturbed problems
[…]
65L11, 68T07, 65T60
1 Introduction
Singularly perturbed problems (SPPs) can be described using differential equations featuring at least one small parameter, popularly known as the “perturbation parameter(s)”. Specifically, when the parameter is very small (close to zero), the solution exhibits thin transition layers, often adjacent to the boundaries of the domain of interest. Consequently, the solution or its derivatives undergo rapid fluctuations within specific domain areas while maintaining smooth behavior outside these regions [1]. These problems are encountered in various physical systems that simultaneously contain both slow and fast varying components in the fields of science and engineering such as fluid dynamics [2, 3, 4], chemical reactors [5], financial modeling [6], gas porous electrodes [7], mathematical biology [8], modeling of semiconductor devices [9], and convective heat transfer problems with high Péclet numbers [10].
The numerical solutions of these parameter-dependent problems based on the discretization approaches are impractical since it is affected by the parameter. Further, a large number of mesh points are required to produce sufficiently precise approximations of the actual solution for various perturbation parameter values. Many methods have been developed to address this challenge, such as parameter-uniform, parameter-robust, or uniformly convergent approaches like Fitted Operator Methods (FOMs) [11, 12, 13] and Fitted Mesh Methods (FMMs) [14, 15]. Moreover, to address the challenges with meshes, various meshfree methods, such as the moving least squares (MLS) method [16], element-free Galerkin approach [17], and the local Petrov–Galerkin method [18], have been developed to solve SPPs numerically [19, 20].
This paper concentrates on one such class of methods, which eliminates not only meshes but also has the ability to handle complex geometry and higher-dimensional problems, namely, physics-informed neural networks (PINNs). A few of the earliest works on solving differential equations utilizing neural networks were presented by Dissanayake et al. [21] and Lagaris et [22]. Their method required a trial solution that satisfies all boundaries and initial conditions. However, this technique often fails if a problem has complex behavior, and finding appropriate trial solutions for such problems becomes challenging. Later, Raissi et al. [23] introduced PINNs, which incorporate initial and boundary conditions into the network’s loss function without considering a trial solution beforehand. Following this work, different types of PINNs, such as variational physics-informed neural networks (VPINN) [24], gradient-enhanced physics-informed neural networks (GPINN) [25], and extended physics-informed neural networks (XPINN) [26], enhanced physics-informed neural networks with Augmented Lagrangian relaxation method (AL-PINNs) [27], and physics-informed extreme learning machines [28, 29], have been developed to enhance the performance of conventional PINN. For more details, see a comprehensive review of PINNs [30, 31] and references therein.
Despite many significant advantages, PINNs encounter challenges in accurately handling equations with rapid oscillations [31]. This difficulty arises because the various components in the loss function have varying convergence rates during training. PINNs need to efficiently represent and learn these features that vary significantly across various scales. While existing theories address issues related to isolated rapid changes, using techniques like domain splitting [32], these approaches can be cumbersome to implement. Furthermore, the authors of [33] introduced an improved physics-informed neural networks algorithm, which involved some modifications from the approach described in [32]. A singularity term was included in the calculation formula to account for the singularity of the equation. Other approaches [34, 35] in which authors sampled dense training points near high-gradient regions to improve convergence. Motivated by the reviewed research and identified shortcomings, this work proposes a novel modified version of PINNs i.e., wavelet-based physics-informed neural networks (W-PINNs), which exploit the ability of wavelets to address these challenges.
The localization properties of wavelets make them beneficial when analyzing problems with localized features, such as initial or boundary layers, in SPPs. Mallat’s work [36] and Daubechies’ contributions [37] lay down the foundational principles of wavelet analysis. Using the Haar wavelet method, Kumar and Deswal [38] have developed two-parameter SPPs with Robin boundary conditions. Further, Kumar et al. designed the Haar wavelet-based approximation scheme for singularly perturbed partial integro-differential equations [39]. Then, Ahsan et al. introduced a wavelet collocation method for solving SPPs with integral boundary conditions [40]. Wavelets excel at capturing nonlinear information within localized regions of the solution. In contrast, neural networks are well-known for learning and modeling global nonlinear relationships. By incorporating wavelets with neural networks, we can potentially leverage the ability of wavelets to extract these localized nonlinearities. This can complement the neural network’s nonlinear learning capability and potentially enhance the overall ability to capture intricate nonlinear patterns, possibly improving the accuracy of the network. The incorporation of wavelets into PINNs has been attempted in [41]. However, this paper focuses on wavelets solely as activation functions, with the overall structure of PINNs remains unchanged from its original introduction. Our findings show that this method also fails to approximate SPPs satisfactorily.
In this study, we use the wavelet family to represent an approximate solution to the differential equation, and the weight of each family member is optimized using the W-PINN. Unlike previous neural network-based approaches for solving SPPs, this method requires no prior information about the nature of the solution, making it practical and easy to implement. Additionally, the proposed neural architecture significantly reduces training time as it does not rely on automatic differentiation for the derivatives involved in the loss function. We use Mexican hat and Gaussian wavelets for implementation, and their performance is compared with conventional PINNs to validate the efficacy. Our approach demonstrates high accuracy across various singularly perturbed ordinary differential equations examined in this research and establishes it as a robust method for these class problems.
The organization of the remaining paper is outlined as follows: Section 2 initiates with the introduction of standard PINNs, followed by a thorough discussion on the design and operational mechanism of W-PINNs. Section 3 presents numerical results for various SPPs to demonstrate the effectiveness and accuracy of the proposed method. Finally, Section 4 serves as the concluding segment, offering a concise summary of the key findings, contributions, and future work along with limitations.
2 Methodology
2.1 Physics-informed neural networks (PINNs)
A neural network is a mathematical model consisting of layers of neurons interconnected via nonlinear operations. These layers include an input layer for initial data, multiple hidden layers for complex computations, and an output layer for predictions. It is widely acknowledged that with enough hidden layers and neurons per layer, a neural network can approximate any function [42]. The values of neurons in each layer depend on the previous layer in the following manner:
[…]
where $\boldsymbol{z}^{k}\in\mathds{R}^{n_{k}}$ is the outcome of $k$-th layer, $\boldsymbol{\omega}^{k}\in\mathds{R}^{n_{k}\times n_{k-1}}$, $\boldsymbol{b}^{k}\in\mathds{R}^{n_{k}}$, and $n_{k}$ is the number of neurons in $k^{th}$ layer. Here, $\boldsymbol{\omega}$ and $\boldsymbol{b}$ denote the weight and bias, the model parameters to be optimized. Moreover, $\sigma$ is a nonlinear mapping known as an activation function. This process is widely known as feed-forward propagation. We define a loss function ($\mathcal{L}$) for network output, which quantifies the disparity between the network’s predictions and the desired outcomes. The neural network is trained using backpropagation [43], a technique wherein optimization algorithms, such as gradient descent, to minimize loss function and fine-tune the model parameters ($\boldsymbol{\omega},\boldsymbol{b}$). Trained parameters capture the underlying structure of the problem in consideration.
Figure 1: A PINN architecture for ODEs.
It is well known that PINNs incorporate the underlying physics of the system by using a governing set of equations along with initial and boundary conditions into the loss function. It measures how much the network’s predicted solution violates the differential equation at several collocation points. Mathematically, it can be written as $\mathcal{L}=\mathcal{L}_{r}+\mathcal{L}_{c}$, where $\mathcal{L}_{r}$ denotes the residual’s mean squared error and $\mathcal{L}_{c}$ denotes the initial and boundary condition’s mean squared error. During optimization, a derivative of the loss function with respect to the model parameters and derivatives of the network’s output with respect to input parameters, exhibited in the loss function are required. These derivatives are effectively computed using “automatic differentiation” [44]. Thus, the objective of PINNs is to identify the optimal parameters that minimize the specified loss function, which encapsulates all physics-related information. Consider a general ordinary differential equation (ODE):
$\begin{cases}\mathscr{L}\big{(}x,u,u^{(1)},u^{(2)},\cdots,u^{(n)}\big{)}=0,~{}~{}x\in(a,b),&\\ \mathscr{B}\big{(}x(a),u(a),u^{(1)}(a),\cdots u^{(n-1)}(a),x(b),u(b),u^{(1)}(b),\cdots u^{(n-1)}(b)\big{)}=0,&\end{cases}$ (2)
where $\mathscr{L}$ denotes a differential operator and $\mathscr{B}$ represents initial and/or boundary conditions. To solve such ODEs, a fully connected schematic PINN architecture is shown in Fig (1).
2.2 Wavelet-based physics-informed neural networks (W-PINNs)
We start with a family of wavelets, whose members are dilates and translates of the mother wavelet. Indeed, the dilated and translated version of the mother wavelet function, $\psi(x)$ for $j,k\in\mathbb{Z}$ can be written as
$\psi_{j,k}(x)=\sqrt{2^{j}}\psi(2^{j}x-k),$
Here $j$ is the exponent of dilation, and $k$ is the translation parameter. In order to define the number of family members, we fix the maximum resolution for the family as $J$. For a given $J$, the dilation exponent, $j$, ranges from 00 to $J-1$, and for a $j,k$ ranges from 00 to $12^{j}-$ 1. This way of constructing a family enables us to cover the whole domain, $[0,1]$ with various resolutions. For instance, $J=2$ will consist of three members: $\psi_{0,0}(x),\psi_{1,0}(x),\psi_{1,1}(x)$.
Figure 2: A W-PINN architecture for second order ODEs.
Figure 3: A W-PINN architecture for coupled second order ODEs.
After constructing a family of wavelets, we approximate the solution of the differential equation as:
$\hat{u}(x)=\sum_{j=0}^{J-1}\sum_{k=0}^{2^{j}-1}c_{j,k}\psi_{j,k}(x)+\mathcal{B}~{}~{}\text{equivalently,}~{}~{}\hat{u}(x)=\sum_{i=0}^{2^{J}-1}c_{i}\psi_{i}+\mathcal{B},$ (3)
where $\mathcal{B}$ is bias, $i=2^{j}+k-1$ and $c_{i}$ denotes the coefficient of the $i^{th}$ member. In this study, we consider two mother wavelets, namely, the Gaussian and the Mexican Hat, which are single and double derivatives of the Gaussian function, respectively. The mathematical expressions for the Mexican hat $(\psi^{M}(x))$, Gaussian $(\psi^{G}(x))$ can be written as
$\psi^{M}(x)=(1-x^{2})e^{-\frac{x^{2}}{2}},~{}~{}~{}~{}\psi^{G}(x)=-xe^{-\frac{x^{2}}{2}}$
Here, we added a bias $\mathcal{B}$ to make a degree of freedom in approximating solutions with non-zero means, as the Mexican Hat and Gaussian are symmetrical wavelets. This enables the model to search for a solution in a space that is more suitable for singularly perturbed problems.
We employ W-PINN to obtain an approximate solution, as described above. W-PINN is composed of two parts: a trainable network that predicts coefficients by forming a fully connected network, and a non-trainable network that provides the final approximated solution and its derivatives. Once we have an approximate solution, we get the loss function of the network using the mean square of residuals and the mean square error of the initial and boundary conditions. The schematic diagrams of W-PINN for second-order ODEs and coupled ODEs are shown in Figure 2 and Figure 3, respectively. W-PINN requires collocation points spanning all over the domain as a training set and a wavelet matrix and its derivative matrices up to the order of the differential equation in consideration. These wavelet matrices serve as the weights of the non-trainable network. A trainable bias is also added to the wavelet matrix $(\mathbf{W})$ in order to incorporate $\mathcal{B}$. For a given set of collocation points $\mathbf{x}=(x_{0},x_{1},\cdots,x_{N})$, the wavelet matrices are constructed as follows:
$\mathbf{W}=\begin{pmatrix}\psi_{0,0}(x_{0})&\cdots&\psi_{j,k}(x_{0})&\cdots&\psi_{J-1,2^{J-1}-1}(x_{0})\\ \psi_{0,0}(x_{1})&\cdots&\psi_{j,k}(x_{1})&\cdots&\psi_{J-1,2^{J-1}-1}(x_{1})\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ \psi_{0,0}(x_{N})&\cdots&\psi_{j,k}(x_{N})&\cdots&\psi_{J-1,2^{J-1}-1}(x_{N})\\ \end{pmatrix},~{}\mathbf{D_{1}W}=\begin{pmatrix}\psi^{\prime}_{0,0}(x_{0})&\cdots&\psi^{\prime}_{j,k}(x_{0})&\cdots&\psi^{\prime}_{J-1,2^{J-1}-1}(x_{0})\\ \psi^{\prime}_{0,0}(x_{1})&\cdots&\psi^{\prime}_{j,k}(x_{1})&\cdots&\psi^{\prime}_{J-1,2^{J-1}-1}(x_{1})\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ \psi^{\prime}_{0,0}(x_{N})&\cdots&\psi^{\prime}_{j,k}(x_{N})&\cdots&\psi^{\prime}_{J-1,2^{J-1}-1}(x_{N})\\ \end{pmatrix},$
where $\psi^{\prime}(x)$ is single derivate of $\psi(x)$ and similarly $\mathbf{D_{2}W}$ is constructed by double derivative of $\psi(x)$.
3 Results and discussions
In this section, we present the results obtained using the proposed W-PINN method over a handful of SPPs and compare them with the performance of conventional PINNs. For a fair comparison, all network parameters are kept the same for both and are tabulated along with the examples. We utilize the Adam optimizer [45], which combines two stochastic gradient descent algorithms: gradient descent with momentum and root mean squared propagation (RMSProp). RMSprop, also known as adaptive gradient descent, adapts the learning rate with iterations. Adam optimizer is memory efficient and is well suited for problems with high-dimensional parameter spaces.
The trainable parameters of networks are initialized using Xavier’s technique or Glorot initialization [46], which, in contrast to random initialization, initializes parameters in a suitable range to achieve faster convergence during training. During backpropagation, it maintains gradients within a reasonable range (not too low or high), thereby preventing the algorithm from getting stuck in poor local minima. Besides this, it can also reduce the need for extensive hyperparameter tuning, particularly the learning rate. We sampled the training set by Sobol sequencing throughout the domain as we need efficient exploration of the entire input space for a single training set. For evaluation purposes, we employ the relative $L_{2}$ metric over uniform samples across the domain.
$\text{Relative}~{}L_{2}~{}\text{error}=\sqrt{\frac{\sum_{i=1}^{M}(u_{i}-\hat{u}_{i})^{2}}{\sum_{i=1}^{M}u_{i}^{2}}},$ (4)
where $M$ represents the total number of testing samples, $u_{i}$ is the exact solution and $\hat{u}_{i}$ is the predicted one for the $i^{th}$ sample. In the case of problems with unknown solutions, we consider the exact solution as a numerical solution obtained from the Scipy solver.
For implementation purposes, the pytorch framework is used. Training is done on the NVIDIA RTX A6000, which has 48GB of GPU memory and 128GB of RAM. Conventional PINN requires roughly four minutes to complete $10,000$ iterations with a network having eight hidden layers, each containing $200$ neurons for $10,000$ collocation points. In contrast, W-PINN requires about one minute, which is significantly less. One of the reasons is that, as mentioned earlier, W-PINN does not require automatic differentiation for derivatives of the network’s output with respect to input parameters. Another reason for faster convergence is that the wavelet space is more suitable for addressing problems with sharp singularities.
As the performance of a neural network is highly sensitive to the initialization of its parameters, we performed each experiment five times and discussed the best of the five here. We begin with a singularly perturbed linear advection-diffusion ODE, then proceed to nonlinear and coupled nonlinear ODEs. Using one of the nonlinear ODEs, we investigate how the network’s hyperparameters affect its performance and also compare it with the approach proposed in [41]. Finally, we tested W-PINN over the FitzHugh-Nagumo model and predator-prey model.
Example 3.1.
Consider the following one-dimension linear advection-diffusion equations [47]:
$\begin{cases}\epsilon\dfrac{d^{2}u}{dx^{2}}+(1+\epsilon)\dfrac{du}{dx}+u=0,~{}x\in(0,1),\\[6.0pt] u(0)=0,~{}u(1)=1.\end{cases}$ (5)
Here, $0<\epsilon\ll 1$. The exact solution to the considered problem is
$u(x)=\dfrac{\exp(-x)-\exp(-\frac{x}{\epsilon})}{\exp(-1)-\exp(-\frac{1}{\epsilon})}.$
Such models, called Friedrich’s boundary layer models, are often used to show how difficult it is to model viscous flow boundary layers [32]. It is impossible for PINNs to capture the singularity that happens when $\epsilon\rightarrow 0$. In [32], improved PINNs are used to solve boundary layer problems with singular perturbation. In which the domain of the problem is conceptually divided into inner and outer regions to improve the application of PINNs. While the method is promising, its computational complexity and implementation challenges limit its practicality. W-PINNs offer significant advantages over this method.
Table 1: (Example 3.1) Network’s parameters.
Parameters Value
Level of resolution ($J_{M}/J_{G}$) 8/10
Number of hidden layers 6
Neurons per layer 100
Learning rate $10^{-4}$
Number of iterations $2\times 10^{4}$
Number of collocation points $10^{3}$
Table 2: (Example 3.1) Relative $L_{2}$-error.
PINN W-PINN
$\epsilon~{}~{}$ Mexican Gaussian
$2^{-1}$ 1.6e-04 5.8e-04 4.1e-04
$2^{-5}$ 3.3e-02 1.9e-04 2.7e-04
$2^{-7}$ 8.4e-01 9.2e-04 5.3e-04
Figure 4: Comparision of PINN with W-PINN for Example 3.1.
The parameters of the neural network for the Example 3.1 are described in Table 2. The results have been presented in Table 2 and Figure 4. Specifically, Table 2 provides the relative $L_{2}$-error and compares the W-PINN with the conventional PINN. This table shows that when $\epsilon=2^{-1}$, there is no singularity near the origin, and PINN approximates the solution well. However, when $\epsilon$ decreases, the exact solution of the problem has a strong singularity near the origin. This means that the layer at the origin becomes very sharp, and PINN is unable to accurately capture the behavior of the solution. W-PINN effectively resolves the singularity, with the approximate solution closely following the exact solution. Similar observations are drawn from Figure 4. It is shown in this figure that for $\epsilon=2^{-7}$, the W-PINN solution precisely matches the exact solution, whereas the PINN solution is far away from the exact one. Therefore, it can be concluded that for the problem, the W-PINN effectively handles the initial singularity effectively.
Example 3.2.
Consider the following highly singularly perturbed non-linear problem [48]:
$\begin{cases}\epsilon\dfrac{d^{2}u}{dt^{2}}+(3+t)\dfrac{du}{dt}+u^{2}-\sin(u)=f,~{}t\in(0,1],\\[6.0pt] u(0)=1,~{}u^{\prime}(0)=1/\epsilon,\end{cases}$ (6)
where f is chosen such that $u(t)=2-\exp(-t/\epsilon)+t^{2}$ is the exact solution.
Table 3: (Example 3.2) Network’s parameters.
Parameters Value
Level of resolution ($J_{M}/J_{G}$) 10/11
Number of hidden layers 8
Neurons per layer 200
Learning rate $10^{-5}$
Number of iterations $2\times 10^{4}$
Number of collocation points $10^{4}$
Table 4: (Example 3.2) Relative $L_{2}$-error.
PINN W-PINN
$\epsilon~{}~{}$ Sigmoid Gaussian Mexican Gaussian
activation activation
$2^{-4}$ 1.2e-04 2.1e-04 1.8e-04 2.3e-04
$2^{-7}$ 8.5e-02 2.2e-02 3.7e-04 1.3e-04
$2^{-10}$ 1.4 0.34 4.5e-03 9.6e-05
Figure 5: Comparision of PINN with W-PINN for Example 3.2.
This example corresponds to a highly singularly perturbed non-linear equation with a known solution. Such an equation is a well-known tool that is frequently exploited to analyze problems in fluid dynamics. for Example 3.2, the parameters of the network are given in Table 4. The results are presented in Table 4 and Figure 5. We compare the W-PINN with the traditional PINN, and the PINN proposed in [41] (using Gaussian wavelet as an activation function) for various $\epsilon$ by considering the relative $L_{2}$-error in Table 4. As the singularity becomes more prominent, Table 4 illustrates that the W-PINN using Gaussian wavelet resolves the singularity to a greater extent.
The error obtained using PINN is very large for $\epsilon=2^{-10}$; in this case, PINNs are unable to detect the singularity, and also PINNs with the Gaussian wavelet as activation function [41] couldn’t resolve singularity satisfactorily, whereas W-PINN provides less error and effectively handles the singularity. One can get corresponding observations from Figure 5, in which solutions using PINN and W-PINN are plotted for various $\epsilon$.
In addition, we investigate the relationship between the hyperparameters of W-PINN and its performance for this example. Figure 7 illustrates the relationship between the resolution level and the relative error. Here, we employ a network with $8$ hidden layers, each containing $200$ neurons, and fix $10,000$ collocation points. It shows that W-PINN gives optimal results for a specific resolution range (not too high nor too low). Lower resolution is incapable of capturing sharp singularities, and a large $J$ makes the loss function too complex to be optimized. The optimal range of the resolution is also dependent on the wavelet family. In this specific case, the W-PINN with Gaussian performed well for $J$, which is between $8$ and $14$, while this range is narrower for the Mexican hat. However, W-PINN performance also depends on the kind of wavelet used for a particular problem.
Figure 7 presents the effect of the number of collocation points on performance. We consider the identical network as previously used, with the Gaussian resolutions ($J_{G}=12$) and the Mexican hat resolutions ($J_{M}=11$). It is evident that the relative $L_{2}$-error decreases as the number of collocation points increases. However, the error decrease is not significant beyond a certain threshold. After a certain number of collocation points, additional points do not come with any new information about the behavior of the problem in consideration.
Figure 8 presents experiments on the trainable part of W-PINN. We employ $10,000$ collocation points for both Gaussian and Mexican, with resolutions of 12 and 11, respectively. It implies that a shallow network is incapable of learning the behaviors of the problem, regardless of the number of neurons per layer. However, a network with a significant number of layers and sufficient neurons per layer performs adequately. It is redundant to employ a large number of layers and neurons, which results in an increase in computational cost rather than accuracy.
Figure 6: Resolution levels $(J)$ versus relative $L_{2}$-error.
Figure 7: Number of collocation points $(N)$ versus relative $L_{2}$-error.
$(a)\hskip 2.84544pt$ Gaussian wavelet
$(b)\hskip 2.84544pt$ Mexican wavelet
Figure 8: Impact of various hidden layers and number of neurons per layer over relative $L_{2}$-error.
Example 3.3.
Consider the following singularly perturbed nonlinear problem with Neumann boundary conditions
$\begin{cases}-\epsilon\dfrac{d^{2}u}{dt^{2}}+u^{5}+3u-1=0,~{}t\in[0,1],\\[6.0pt] u^{\prime}(0)=\sin(0.5),~{}u^{\prime}(1)=\exp(-0.7).\end{cases}$ (7)
This example corresponds to a singularly perturbed nonlinear problem with Neumann boundary conditions. This problem possesses boundary layer singularity at both ends. The exact solution to this problem is unknown, so we obtained a numerical solution using Scipy solver and treated it as an exact solution.
Table 6 represents the parameters of the network for the example 3.3. The results have been displayed in Table 6 and Figure 9. Specifically, the table 6 compares the performance of the W-PINN and traditional PINN in terms of relative $L_{2}$-error across various $\epsilon$ values. By analyzing this table, it is evident that as the singularity becomes more prominent, our W-PINN effectively addresses the singularity to a greater degree compared to the PINN. Furthermore, Figure 9 is plotted for various $\epsilon$ values. This figure illustrates that as $\epsilon$ decreases, PINN encounters difficulties in capturing these singularities, as it is unable to incorporate them into its solutions. However, W-PINN efficiently adapts these singularities.
Table 5: (Example 3.3) Network’s parameters.
Parameters Value
Level of resolution ($J_{M}/J_{G}$) 9/12
Number of hidden layers 8
Neurons per layer 200
Learning rate $10^{-4}$
Number of iterations $3\times 10^{4}$
Number of collocation points $10^{4}$
Table 6: (Example 3.3) Relative $L_{2}$-error.
PINN W-PINN
$\epsilon~{}~{}$ Mexican Gaussian
$2^{-4}$ 8.6e-05 1.4e-04 4.7e-05
$2^{-7}$ 1.8e-03 8.3e-05 8.9e-06
$2^{-10}$ 5.3e-03 9.9e-05 6.5e-06
Figure 9: Comparision of PINN with W-PINN for Example 3.3.
Example 3.4.
Consider the following singularly perturbed non-linear coupled problem [49]
$\begin{cases}\epsilon_{1}\dfrac{du_{1}}{dx}+3u_{1}-\dfrac{1}{4}\exp(-u_{1}^{2})-u_{2}-x^{2}+1=0,~{}x\in(0,1],\\[6.0pt] \epsilon_{2}\dfrac{du_{2}}{dx}+4u_{2}-\cos(u_{2})-u_{1}-\exp(x)=0,~{}x\in(0,1],\\[6.0pt] u_{1}(0)=1,~{}u_{2}(0)=1.\end{cases}$ (8)
This example represents a system of non-linear singularly perturbed differential equations and has an unknown solution. In particular, we investigate three cases: The case-1 occurs when there is no singularity in the solution $u_{1}$, but the solution $u_{2}$ contains an abrupt singularity, i.e., $\epsilon_{1}=1$ and $\epsilon_{2}=2^{-10}$. The case-2 is the other way around; such a model allows us to study how fast and slow processes interact and influence the overall system behavior. The case-3 refers to singularity in solutions $u_{1}$ and $u_{2}$, i.e., $\epsilon_{1}=\epsilon_{2}=2^{-10}$. This case represents scenarios where both processes occur simultaneously and rapidly impact the overall behavior of the system.
In our experiment for the Example 3.4, we use the parameters for the network in Table 8. Table 8 and Figure 10 represent the solutions. Specifically, we present the relative $L_{2}$-error with the W-PINN and the traditional PINN for various $\epsilon_{1}$ and $\epsilon_{2}$ in Table 8. From this table, we also observe that the error is much less for W-PINN compared to PINN. The similar performance difference between PINN and W-PINN is evident in Figure 10.
Table 7: (Example 3.4) Network’s parameters.
Parameters Value
Level of resolution ($J_{M}/J_{G}$) 10/12
Number of hidden layers 8
Neurons per layer 200
Learning rate $10^{-5}$
Number of iterations $2\times 10^{4}$
Number of collocation points $10^{4}$
Table 8: (Example 3.4) Relative $L_{2}$-error.
PINN W-PINN
$\epsilon_{1}$
$\epsilon_{2}$ Mexican Gaussian
111~{}~{}~{}~{}1 $u_{1}$ 9.8e-03 2.8e-04 1.1e-04
$2^{-10}$ $u_{2}$ 1.2e-02 5.2e-03 6.7e-04
$2^{-10}$ $u_{1}$ 7.1e-01 8.1e-03 5.7e-03
111~{}~{}~{}~{}1 $u_{2}$ 3.2e-02 5.1e-04 1.8e-04
$2^{-10}$ $u_{1}$ 5.4e-01 7.1e-03 6.8e-03
$2^{-10}$ $u_{2}$ 4.9e-02 1.1e-03 8.9e-04
Figure 10: Comparision of PINN with W-PINN for Example 3.4.
Table 9: (Example 3.5) Network’s parameters.
Parameters Value
Level of resolution ($J_{M}/J_{G}$) 11/12
Number of hidden layers 10
Neurons per layer 300
Learning rate $10^{-5}$
Number of iterations $4\times 10^{4}$
Number of collocation points $10^{4}$
Table 10: (Example 3.5) Relative $L_{2}$-error.
PINN W-PINN
$\tau~{}~{}$ Mexican Gaussian
$2^{-10}$ $v$ 5.9e-04 4.4e-03 9.2e-05
$w$   8.1e-02 8.2e-03 5.7e-04
Figure 11: Comparision of PINN with W-PINN for Example 3.5, using $\tau=2^{-10}$.
Example 3.5.
FitzHugh-Nagumo (FHN) model [50]:
The FitzHugh-Nagumo (FHN) model is a simplified version of the Hodgkin-Huxley model, which simulates the dynamics of spiking neurons. It captures the core aspects of how a neuron generates electrical signals (spikes) in response to stimulation (an external current). The mathematical form of the FHN dynamical model is defined as follows:
$\begin{cases}\dfrac{dv}{dt}-v+\dfrac{v^{3}}{3}+w-RI=0,\\[6.0pt] \tau\dfrac{dw}{dt}-v+bw+a=0,\\ \end{cases}$ (9)
where $v$ denotes the membrane voltage of the neuron, $w$ is the recovery variable, reflecting the activation state of ion channels, $I$ represents the external current applied to the neuron, and $R$ is the resistance across the neuron. In addition, $a$ and $b$ are scaling parameters, and $\tau$ denotes the time constant for the recovery variable ($w$), which acts as a singularly perturbed parameter for this model. In particular, $a=b=0$, the FHN model describes the Van der Pol oscillator, which describes self-sustaining oscillations in many systems, such as heartbeats, economies, and electronic circuits.
For the numerical computations, we set: $a=1,~{}b=1,~{}I=0.1,~{}R=1,$ and $\tau=2^{-10}$, $v(0)=0.5$ and $w(0)=0.1$. The dynamics of the system are computed till $t=1$. Table 10 provides the parameters for this example 3.5. The results are presented in Table 10 and Figure 11. In particular, the relative $L_{2}$-error with the W-PINN and the traditional PINN is shown in Table 10. We present the W-PINN approximations of $v$ and $w$ in Figure 11. This figure is evident that an initial layer is present in the solution profile of $w$. Further, the figure illustrates that PINN fails to capture the sharp singularity of the solution and generates spurious oscillations near the singularity, whereas W-PINN effectively resolves the singularity.
Table 11: (Example 3.6) Network’s parameters.
Parameters Value
Level of resolution ($J_{M}/J_{G}$) 11/13
Number of hidden layers 8
Neurons per layer 200
Learning rate $10^{-5}$
Number of iterations $3\times 10^{4}$
Number of collocation points $10^{4}$
Table 12: (Example 3.6) Relative $L_{2}$-error.
PINN W-PINN
$\epsilon~{}~{}~{}~{}~{}~{}~{}~{}$ Sigmoid Gaussian Mexican Gaussian
activation activation
$2^{-8}$  $x$ 0.342 5.8e-03 4.6e-03 6.3e-04
$y$ 0.997 0.991 1.8e-03 6.7e-04
Figure 12: Comparision of PINN with W-PINN for Example 3.6, using $\epsilon=2^{-8}$.
Example 3.6.
Predator-prey interaction model [51]:
A predator-prey system is an interconnected population of two species in which one species, the predator, actively hunts and consumes the other species, the prey. A singular perturbation argument detects limit cycle behavior and describes other dynamical situations observed in the predator-prey interaction. The Predator-prey interaction model is mathematically defined as follows:
$\begin{cases}\dfrac{dx}{dt}=kxy-\phi_{1}(x),\\[6.0pt] \dfrac{dy}{dt}=\phi_{2}(y)-\dfrac{1}{\epsilon}xy,\\ \end{cases}$ (10)
where $x(t)\geq 0$ denotes the predator density, $y(t)\geq 0$ represents the prey density, $\phi_{1}(x)$ denotes the natural death rate of the predator population, and $\phi_{2}(y)$ denotes the change in prey population due to natural growth. Here, we consider $\phi_{1}(x)=dx$, which denotes the exponential decay rate of the predator population due to natural death. Moreover, $\phi_{2}(y)=ry\Bigl{(}1-\dfrac{y}{m}\Bigr{)}$, with $r>0$, $m>0$ is a logistic growth rate for the prey population. The parameter $k$ is referred to as the conversion rate.
For the numerical computations, we set $k=0.5,~{}d=1,~{}r=1,~{}m=0.1,$ and $\epsilon=2^{-8}$ (singular perturbation parameter), let the initial density of both predator and prey be same, $x(0)=1$ and $y(0)=1$. The dynamics of the system are computed till $t=1$.
We consider parameters for this problem 3.6 that are presented in Table 12. The results for this example are discussed in Table 12 and Figure 12. Specifically, Table 12 displays a comparison between the W-PINN, the traditional PINN, and the PINN described in [41] that employs a Gaussian wavelet as an activation function for $\epsilon=2^{-8}$. It is evident from this table that the W-PINN with a Gaussian wavelet addresses the singularity more effectively. Further, the large errors in the $x$ and $y$ solutions obtained with the PINN using a sigmoid activation function highlight its failure to detect the singularity. While the PINN with the Gaussian wavelet as an activation function accurately approximates the solution $x$ without any singularity, it fails to satisfactorily resolve the singularity in the solution $y$. On the other hand, the W-PINN exhibits lower error and effectively handles the singularity in the $y$ solutions. These observations can also be drawn from Figure 12, in which solutions using PINN and W-PINN are plotted for $\epsilon=2^{-8}$.
4 Conclusion
This paper successfully demonstrated the advantages of W-PINNs over traditional PINNs and PINNs with wavelet activation functions for addressing singularly perturbed problems. The wavelet theory served as the foundation for the development of W-PINN, which combined the learning efficiency of PINN with the localization properties of wavelets (both in scale and space) to capture nonlinear information within localized regions. A limitation of the proposed method is that it will perform adequately when the resolution of the wavelet family is chosen over a specific small interval. A natural extension of this work involves applying it to singularly perturbed partial differential equations. W-PINNs may also prove effective for other problems that exhibit singular behavior, such as fractional differential models, which inherently possess such characteristics.
CRediT authorship contribution statement
Himanshu Pandey: Conceptualization, Methodology, Programming, Formal analysis, Writing - original draft. Anshima Singh: Conceptualization, Investigation, Methodology, Formal analysis, Writing - original draft. Ratikanta Behera: Funding acquisition, Supervision, Writing - review & editing.
Declaration of competing interest
The authors declare that there are no conflicts of interest.
Data Availability
No data was used for the research described in the article.
Acknowledgments
Ratikanta Behera acknowledges the support of the Science and Engineering Research Board (SERB), Department of Science and Technology, India, under Grant No. EEQ/2022/001065.